Prove the trigonometric identity: sin(x + y) sin(x - y) = sin² x – sin? y. Which identity is used to prove it true?
Answer #1

We Sin (22+4) Sin (C-4) = Sin²x - Sin?y know that, Sin (A+B) = (SinA GAB & Cox A Sin B) and Sin (A-B) = (SinA GAB Coxa Sin B)